find equation of a line when slope is -2 and passing through (1,2)
Found 2 solutions by jim_thompson5910, Sarah900:
Answer by jim_thompson5910(35256)


y = mx+b

2 = -2(1)+b

2 = -2+b

2+2 = b

4 = b

b = 4

So the equation of the line is y = -2x+4


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

Answer by Sarah900(12)

To solve this you are going to use the formula y – y1 = m(x – x1). From there plug in what you know. m is the slope which for this problem is -2. x1 is the x coordinate of the given point. In this case its 1. And y1 is the y coordinate of the given point. In this case its 2.
y-2 = -2(x – 1)
Now simply manipulate the problem into slope intercept form
y-2=-2x+2
y=-2x+4
To check our answer you can graph it. If the line goes through the given point, then your answer is correct.